### v1

# On the metric property of quantum Wasserstein divergences

Abstract.
Quantum Wasserstein divergences are modified versions of quantum Wasserstein distances defined by channels, and they are conjectured to be genuine metrics on quantum state spaces by De Palma and Trevisan. We prove triangle inequality for quantum Wasserstein divergences for any finite-dimensional quantum system and any quadratic cost operator under the assumption that a particular state involved is pure. We also provide strong numerical evidence suggesting that the triangle inequality holds in general, for an arbitrary choice of states.
Key words and phrases:
quantum optimal transport, metric property2020 Mathematics Subject Classification:
Primary: 49Q22; 81P16. Secondary: 81Q10.Contents
- 1 Introduction
- 2 Triangle inequality for quantum Wasserstein divergences — the proof of Theorem 1
- 3 Numerical evidence for the triangle inequality for generic triplets of states
- 4 Steps towards an analytic proof of the triangle inequality in full generality
- 5 Appendix: the code used for the numerical study
1. Introduction
1.1. Motivation and main result
Although the classical problem of transporting mass in an optimal way was formulated in the 18th century by Monge, the theory of classical optimal transport (OT) became one of the central topics of analysis only in the last few decades with intimate links to mathematical physics [15, 33, 35], PDE theory [1, 22, 36], and probability [2, 3, 31]. Not to mention the countless applications in artificial intelligence, image processing, and many other fields of applied sciences. See e.g. [21, 38, 37] and the references therein.
The quantum counterpart of classical OT is just emerging in these years. As always, the correspondence between the classical world and the quantum world is not one-to-one. Non-commutative optimal transport is a flourishing research field these days with several different promising approaches such as that of Biane and Voiculescu [4], Carlen, Maas, Datta, and Rouzé [11, 12, 10, 14, 13], Caglioti, Golse, Mouhot, and Paul [8, 9, 25, 27, 26, 28], De Palma and Trevisan [16, 17], Życzkowski and his collaborators [23, 42, 43, 5], and Duvenhage [19, 20]. Separable quantum Wasserstein distances have also been introduced recently [40]. From our viewpoint, the most relevant approach is the one of De Palma and Trevisan involving quantum channels, which is closely related to the quantum optimal transport concept of Caglioti, Golse, Mouhot, and Paul based on quantum couplings.
It is a common feature of both the channel-based and the coupling-based quantum optimal transport distances that they are not genuine metrics — in particular, states may have a positive distance from themselves. This phenomenon is natural on the one hand, considering the nature of quantum mechanics, but has counter-intuitive consequences on the other hand. One of these consequences is that there exist non-surjective and even non-injective quantum Wasserstein isometries (i.e., distance preserving maps) on the compact space of states of a finite-level quantum system [24] — none of these could possibly happen in a genuine metric setting.
As a response to this phenomenon, De Palma and Trevisan introduced quantum Wasserstein divergences [18] which are appropriately modified quantum Wasserstein distances — see (14) for a precise definition. They conjectured that quantum Wasserstein divergences are genuine metrics on quantum state spaces. This paper is devoted to the question whether the triangle inequality holds for these divergences. We formulate our main result in an informal way below — see Theorem 1 for the precise statement.
Main result.
For every finite-level quantum system and for every finite collection of observable quantities, the corresponding quadratic quantum Wasserstein divergence satisfies
for any triplet of states assuming that is pure or both and are pure. Moreover, numerical results strongly indicate that (1) holds for any triplet of states without any further assumptions.
This paper is organized as follows. In the next subsection (Subsec. 1.2) we introduce all the necessary notions and notation. Section 2 is dedicated to the analytical proof of Theorem 1, which is a precise formulation of the above main result. In Section 3 we present numerical evidence suggesting that the conclusion of Theorem 1 holds without any extra assumptions on the states involved. Encouraged by the numerics presented in Section 3, we make steps in the direction of an analytic proof of the triangle inequality in full generality — this material is presented in Section 4.
Finally, we remark the analogy of our present problem with the problem of the metric property of the quantum Jensen-Shannon divergence. In that case, the triangle inequality was proved in 2008 for pure states analytically and for mixed states numerically [34] (see also [6]). A decade later, the problem was settled by an analytical proof in full generality — see [41] and [39].
1.2. Basic notions, notation
The classical OT problem is to arrange the transportation of goods from producers to consumers in an optimal way, given the distribution of production and consumption (described by probability measures and ), and the cost of transporting a unit of goods from to Accordingly, a transport plan is modeled by a probability distribution on the product of the initial and the target space, where is the amount of goods to be transferred from to and hence the marginals of are and So the optimal transport cost is the infimum of a convex optimization problem with linear loss function:
where denotes the th marginal of and is the initial and is the target space. Under mild continuity assumptions on — lower semi-continuity does the job — the infimum in (2) is actually a minimum as it is realized by a transport plan. These minimizing plans are called optimal transport plans.
One has great freedom in choosing the cost function However, the quadratic cost which is simply the square of the distance, plays a distinguished role. Its importance comes mainly from fluid dynamics and the dynamical theory of optimal transportation — the exponent is distinguished by the fact that the kinetic energy is proportional to the square of the velocity. Accordingly, the relevance of quadratic Wasserstein spaces has grown dramatically in recent decades due to their close connection with PDE theory and gradient flows. Recall that if is complete and separable metric space, then the classical quadratic Wasserstein space is the collection of those probability measures on the Borel -algebra that satisfy for some , endowed with the quadratic Wasserstein distance
where the infimum runs over all couplings of and
In classical mechanics, the state of a particle moving in is described by a probability measure on the phase space which is the collection of all possible values of the position and momentum variables In this concrete setting, the quadratic Wasserstein distance (3) of the classical states is given by
where and — here denotes the th marginal of Recall that Wasserstein distances admit a picturesque probabilistic interpretation as they are defined by optimization over couplings of probability measures. Let us stick to the concrete case of -Wasserstein distances between states of classical mechanical systems. In this case, the probabilistic version of (4) reads as follows:
In the above formula, the random variables and represent the position and the momentum of the th particle (). That is, we minimize the sum of the expected squared differences between the positions of the two particles and the momenta of them.
In quantum mechanics, the state of a particle moving in is described by a wave function of unit norm, or more generally, by a normalized, positive, trace-class operator on Measurable physical quantities correspond to (possibly unbounded) self-adjoint operators on The spectrum of such an operator is precisely the collection of all possible outcomes of a quantum measurement. In the sequel we denote by the real linear space of self-adjoint operators on , and stands for the set of states, that is, the set of positive operators on with unit trace. When measuring an observable quantity on a quantum system being in the state the probability of the outcome lying in an interval is where is the spectral measure of Consequently, a quantum state encapsulates a bunch of classical probability distributions, each corresponding to a physical quantity we are interested in. As the conceptual challenges of quantum optimal transport are present in the finite-dimensional setting, we assume from now on that
Let us single out a few observable quantities we are interested in, and let denote the random variable obtained by measuring in the initial state and let denote the random variable obtained by measuring in the final state According to (5), the squared OT distance of the quantum states should read as
where the infimum is taken over all possible couplings of the quantum states and According to the convention introduced by De Palma and Trevisan [17], the set of all couplings of the quantum states is denoted by and is given by
That is, a coupling of and is a state on such that
and
for all Note the clear analogy of the above definition of quantum couplings with the definition of classical couplings that can be rephrased as follows: is a coupling of and if
and
for every continuous and bounded function defined on We remark that is never empty, because the trivial coupling belongs to .
Note that the definition of couplings (7) proposed by De Palma and Trevisan [17] is different from the definition proposed by Golse, Mouhot, Paul [25] in the sense that it involves the dual Hilbert space and hence the transpose operation. For a clarification of this difference, see Remark 1 in [17]. For more detail on the latter concept of quantum couplings, the interested reader should consult [8, 9, 25, 26, 29, 30, 28].
Let us jump back to (6) and note that if the state of our composite quantum system is described by then
Therefore, in view of (6), (7), and (9), the quadratic Wasserstein distance of and with respect to the measurable quantities is given by
There is a correspondence between quantum couplings of the states (in the sense of (7)), and quantum channels (that is, completely positive, trace-preserving linear maps) defined on the support of that map to . This one-by-one relation is constructive and explicit, see Subsection 2.2. and especially Proposition 2 of [17] for further details. A consequence of this explicit relation — which we will heavily use in the sequel — is that the cost of a coupling can be expressed in terms of the corresponding channel as follows:
See Subsection 5.2. of [18] for the justification of the above relation. Consequently, the quadratic quantum Wasserstein distances are obtained by an optimization over quantum channels:
An important feature of the quadratic Wasserstein distances is that the distance of a state from itself (which may be positive) is always realized by the identity channel — see [17, Corollary 1]. In addition to that, the following concavity-like result is true for any choice of and any :
Indeed, (13) is an easy consequence of Theorem 1 and Corollary 1 of [17]. The following conjecture was popularized by Dario Trevisan and Giacomo De Palma in the fall of 2022 — see also [18].
Conjecture 1 (De Palma-Trevisan, [18]).
A modified version of the quantum optimal transport distance (10) defined by
is a true metric for all cost operator of the form
— up to some non-degeneracy assumptions on the ’s generating to ensure the definiteness of that is, that only if
The fact that the self-distance is realized by the identity channel for every [17, Corollary 1] and (12) together imply that
We denote by the set of rank-one ortho-projections on that is, the set of pure states.
Remark 1.
If either or is pure then the only coupling is the tensor product and the only corresponding channel is , with
2. Triangle inequality for quantum Wasserstein divergences — the proof of Theorem 1
Having introduced all the necessary notions and notation, we are in the position to state and prove our main result.
Theorem 1.
Proof.
If is pure or both and are pure, then by by Remark 1 and (16) we have
and
By relaxing the infimum in the definition of the quantum Wasserstein divergence to the tensor product coupling, we get
By (13) the quantum Wasserstein divergence is a non-negative real number and hence eq. (17) is equivalent to
If and are self-adjoint, and , then by the Cauchy-Schwarz inequality for the Hilbert-Schmidt inner product on and , and hence we get the following upper bound for the right-hand side (RHS) of (21):
Now a Cauchy-Schwartz for the Euclidean space tells us that
where we used again the Cauchy-Schwarz inequality for the Hilbert-Schmidt inner product for This completes the proof of (21) and hence that of the Theorem. ∎
3. Numerical evidence for the triangle inequality for generic triplets of states
The quantum optimal transport problem is a semidefinite programming task, and we used Wolfram Mathematica [32] to perform numerical simulations. The data generated during our experiments (Mathematica notebooks, their pdf images, and the raw data exported from the notebooks) is available online, see [7].
In the quantum bit () case, we ran the following experiment. We chose four pairs of random states and four triples of random self-adjoint operators: and The random states are normalized Wishart matrices: they are of the form where is a random matrix with i.i.d. complex standard Gaussian entries. The random self-adjoint operators are defined similarly: they are of the form where the elements of are i.i.d. complex Gaussians. Then we let run on the following lattice within the state space:
We computed the minimal gap between the two sides of the triangle inequality
for every We found the following elements long list of minimal gaps:
Then we turned to the case of higher dimensions In each dimension, we generated triples of random states
which are i.i.d. normalised Wishart matrices, and triples of random self-adjoint matrices Then we computed the minimal gap
We found that and
That is, we found strong numerical evidence indicating that the conclusion of Theorem 1 holds in full generality, without any additional assumption on the states involved.
These minima are convincing but do not tell too much about how the gap depends on the states involved. Therefore, we worked out the following illustrative examples.
In the first example, we considered and chose the deterministic states and We singled out the section of the Bloch ball, that is, we took ’s of the form
where The set of self-adjoint matrices generating the quadratic cost operator is chosen to be We considered the gap
where and plotted it as a function of and in Figure 1.
The second example deals with We chose
and
We considered ’s of the form
and the cost governed by the collection of all possible tensor products of Pauli matrices:
Figure 2 shows the plot of the gap
The following Figure 3 concerns the qubit case again, but now the states and are random (independent normalized Wishart) densities, the quadratic cost operator is generated by random self-adjoint operators, and runs over the section that is, where
Finally, Figure 4 shows the behaviour of the gap in the case, with random states and and random quadratic cost. The third state runs over the region described in (26).
4. Steps towards an analytic proof of the triangle inequality in full generality
The positive numerical results presented in Section 3 encouraged us to take steps towards an analytical proof of the general case. In Subsection 4.1 we study the special case when and the cost operator is the symmetric cost. As we will see, it is vital to get useful lower bounds for quantum Wasserstein distances. One way to get these lower bounds is to find (not necessarily optimal) solutions of the dual optimization problem — see Proposition 2. Another way of getting lower bounds is described in Subsection 4.2. The starting point of this method is the observation that the summands of the quadratic cost operators are gaps between the arithmetic and geometric means of certain operators.
4.1. Triangle inequality for qubits — the case of the symmetric cost
In the special case elements of can be represented by vectors using the Bloch representation. The Bloch vector of a state is defined by
where the ’s are the Pauli operators
The positivity condition is equivalent to the Euclidean length of being at most
In this section we stick to and we choose the cost operator (15) to be
where the “s” in the subscript of refers to the symmetry of meaning that it involves all the Pauli matrices. The induced quantum Wasserstein distance (10) is denoted by and the corresponding modified quantum Wasserstein distance, or Wasserstein divergence (14) is denoted by
We aim to prove the triangle inequality
for a reasonably large class of states
As we have seen, (29) is equivalent to
By the very definition of the Wasserstein divergence (14), the above (30) can be written as
Our strategy to prove (31) is to give a lower bound of the left-hand side of (31) and an upper bound of the right side of (31) such that The following statement will turn out to be useful in deriving such lower and upper bounds.
Proposition 2.
We have the lower bound
for any where is the Bloch vector of and denotes the (Euclidean) norm.
Proof.
The first step is to prove that
for any satisfying A crucial observation of [24] is that the symmetric cost operator defined in (28) is unitary invariant, that is,
for any Therefore, the spectral resolution of which is computed in [24] in the computational basis, is valid in any orthonormal basis.
Let be the eigenbasis of and let be the corresponding dual basis. Then we have
where
Now let us note that always annihilates Indeed,
And on the subspace orthogonal to it is enough to guarantee that
which is clear as both and is bounded from above by So we have proved (33) and now we use it to get useful lower bounds on quantum Wasserstein distances by appropriate choices of The lower bound on the symetric cost (33) implies that for any we have
We aim for the highest possible lower bound on so let us choose as follows:
Moreover, we have the following explicit formula for the self-distance
Proposition 3.
We have
where denotes the Euclidean () norm.
Proof.
The purification of realizes the self-distance: The spectral decomposition of is
and it follows from the spectral resolution of that
where Therefore,
With (32) and (37) in hand, we can give the following upper bound on the right-hand side of (31):
where was bounded by the cost of the independent coupling.
The lower bound on the left hand side of (31) that we can obtain by Proposition 2 and Proposition 3 reads as follows:
We summarize the above computations in the following corollary.
Corollary 4.
Let us choose such that their Bloch vectors satisfy
Then the quantum Wasserstein divergence corresponding to the symmetric cost operator satisfies the triangle inequality
Remark 2.
There are various ways to obtain easy-to-check examples of states that satisfy the assumption of Corollary 4. One way is to fix that is, In this case, simple -variable calculus shows that (38) is satisfied whenever both and are at least no matter what the angle between and is. Moreover, a numerical test shows if we choose a random triplet of states (the states are chosen independently according to the uniform distribution on the Bloch ball), it will satisfy (38) with high probability (around ). The details of this numerics is presented in the subfolder ”Corollary-4” of [7]. Note, however, that there are states that do not satisfy (38): the easiest example is
Remark 3.
Note that the estimate (33) may be sharp: the numerics tells us that it gives the precise transport cost in the special cases
4.2. Lower estimation of the cost function
Assuming that for every we can write the cost operator as the difference of an arithmetic and a geometric mean as follows
Introducing the function
the cost operator can be written as
is an operator convex function on and its tangent line at is given by
By the convexity of we have
and
We can use this estimation to give a lower bound for the cost function in the state .
for all , where and
The function
has maximum at . Since
we deduce that the cost function can be estimated from below by the Hellinger distance of and that is,
5. Appendix: the code used for the numerical study
Here we present the Wolfram Mathematica [32] code that we used to obtain the numerical results discussed in Section 3.
RandSelfadjMatrix[l_] := Module[{A, i}, A = RandomVariate[NormalDistribution[], {l, l}] + I * RandomVariate[NormalDistribution[], {l, l}]; A = A + ConjugateTranspose[A]; A = Chop[A]; A] RandPositive[l_, r_] := Module[{A}, A = RandomVariate[NormalDistribution[], {l, r}] + I * RandomVariate[NormalDistribution[], {l, r}]; A = A.ConjugateTranspose[A]; A] RandState[l_, r_] := Module[{A}, A = RandPositive[l, r]; A = A / Tr[A]; A] CostfromObservable[Observables_, transpose_] := Sum[ MatrixPower[KroneckerProduct[Observables[[k]], IdentityMatrix[Length[Observables[[1]]]]] - KroneckerProduct[IdentityMatrix[Length[Observables[[1]]]], If[transpose,Transpose[Observables[[k]]], Observables[[k]]]], 2], {k, 1, Length[Observables]}] SaMCNB[dim_] := SaMCNB[dim] = Flatten[{{Table[SparseArray[{{k, k} -> 1, {dim, dim} -> 0}], {k, 1, dim}]}, Table[Table[SparseArray[{{k, m} -> Sqrt[2] / 2, {m, k} -> Sqrt[2] / 2, {dim, dim} -> 0}], {m, k + 1, dim}], {k, 1, dim - 1}], Table[Table[SparseArray[{{k, m} -> -I * Sqrt[2] / 2, {m, k} -> I * Sqrt[2] / 2, {dim, dim} -> 0}], {m, k + 1, dim}], {k, 1, dim - 1}]}, 2]; SaMPB[2] = {{{1, 0}, {0, 1}}, {{0, 1}, {1, 0}}, {{0, -I}, {I, 0}}, {{1, 0}, {0, -1}}}; SaMPB[dim_] := SaMPB[dim] = Flatten[Table[ KroneckerProduct[SaMPB[dim / 2][[j]], SaMPB[2][[k]]], {j, 1, dim^2 / 4}, {k, 1, 4}], 1]; SaMPNB[dim_] := SaMPNB[dim] = SaMPB[dim] / Sqrt[dim]; QOT[rho_, omega_, C_, dual_, transpose_] := Module[{dim, sol, Pi, x, y, X, Y}, dim = Length[rho]; If[ dual, {x = Table[Symbol["x" <> ToString[n]], {n, dim^2}]; y = Table[Symbol["y" <> ToString[n]], {n, dim^2}]; sol = SemidefiniteOptimization[-Tr[(y.SaMCNB[dim]).omega + (x.SaMCNB[dim]).rho], {VectorGreaterEqual[{C - KroneckerProduct[(y.SaMCNB[dim]), IdentityMatrix[dim]] - KroneckerProduct[IdentityMatrix[dim], If[transpose, Transpose[ (x.SaMCNB[dim])], x.SaMCNB[dim]]], 0}, {"SemidefiniteCone", dim^2}]}, Flatten[{x, y}]]; X = Chop[sol[[1 ;; dim^2, 2]].SaMCNB[dim], 10^-3]; Y = Chop[sol[[dim^2 + 1 ;; 2 * dim^2, 2]].SaMCNB[dim], 10^-3]; Chop[Sqrt[Tr[X.omega + Y.rho]], 10^-3], X // MatrixForm, Y // MatrixForm}, {x = Table[Symbol["x" <> ToString[n]], {n, dim^4}]; sol = SemidefiniteOptimization[ Chop[Simplify[Tr[(x.SaMCNB[dim^2]).C]], 10^-3], {ResourceFunction["MatrixPartialTrace"] [x.SaMCNB[dim^2], 2, {dim, dim}] -> omega, ResourceFunction["MatrixPartialTrace"] [x.SaMCNB[dim^2], 1, {dim, dim}] -> If[transpose, Transpose[rho], rho], VectorGreaterEqual[{x.SaMCNB[dim^2], 0}, {"SemidefiniteCone", dim^2}] },x]; Pi = Chop[Sum[sol[[n, 2]]SaMCNB[dim^2][[n]], {n, 1, dim^4}], 10^-3]; Chop[Sqrt[Tr[C.Pi]], 10^-3], Pi // MatrixForm}]] ModQOT [rho_, omega_, cost_, dual_, transpose_] := Sqrt[QOT[rho, omega, cost, dual, transpose][[1]]^2 - (QOT[rho, rho, cost, dual, transpose][[1]]^2 + QOT[omega, omega, cost, dual, transpose][[1]]^2) / 2] TriIneq[rho_, omega_, tau_, C_, dual_, transpose_] := ModQOT[rho, omega, C, dual, transpose] + ModQOT[omega, tau, C, dual, transpose] - ModQOT[rho, tau, C, dual, transpose]

### v2

@@
 Quantum Wasserstein divergences are modified versions of quantum Wasserstein distances defined by channels, and they are conjectured to be genuine metrics on quantum state spaces by De Palma and Trevisan.
-We prove triangle inequality for quantum Wasserstein divergences for any finite-dimensional quantum system and any quadratic cost operator under the assumption that a particular state involved is pure.
+We prove triangle inequality for quantum Wasserstein divergences for every quantum system described by a separable Hilbert space and any quadratic cost operator under the assumption that a particular state involved is pure, and all the states have finite energy.
 We also provide strong numerical evidence suggesting that the triangle inequality holds in general, for an arbitrary choice of states.
@@
 - 4 Steps towards an analytic proof of the triangle inequality in full generality
-- 5 Appendix: the code used for the numerical study
+- 5 Applications
+- 6 Appendix: the code used for the numerical study
 1. Introduction
@@
 Main result.
-For every finite-level quantum system and for every finite collection of observable quantities, the corresponding quadratic quantum Wasserstein divergence satisfies
-for any triplet of states assuming that is pure or both and are pure. Moreover, numerical results strongly indicate that (1) holds for any triplet of states without any further assumptions.
+For every quantum system described by a separable Hilbert space and for every finite collection of observable quantities, the corresponding quadratic quantum Wasserstein divergence satisfies
+for any triplet of states assuming that is pure or both and are pure, and all the states involved are of finite energy.
+Moreover, numerical results strongly indicate that (1) holds for any triplet of states without any further assumptions.
 This paper is organized as follows.
@@
 In quantum mechanics, the state of a particle moving in is described by a wave function of unit norm, or more generally, by a normalized, positive, trace-class operator on Measurable physical quantities correspond to (possibly unbounded) self-adjoint operators on The spectrum of such an operator is precisely the collection of all possible outcomes of a quantum measurement.
-In the sequel we denote by the real linear space of self-adjoint operators on , and stands for the set of states, that is, the set of positive operators on with unit trace.
-When measuring an observable quantity on a quantum system being in the state the probability of the outcome lying in an interval is where is the spectral measure of Consequently, a quantum state encapsulates a bunch of classical probability distributions, each corresponding to a physical quantity we are interested in.
-As the conceptual challenges of quantum optimal transport are present in the finite-dimensional setting, we assume from now on that
+In the sequel, we denote by the set of self-adjoint but not necessarily bounded operators on , and stands for the set of states, that is, the set of positive trace-class operators on with unit trace.
+The space of all bounded operators on is denoted by and we recall that the collection of trace-class operators on is denoted by and defied by Similarly, stands for the set of Hilbert-Schmidt operators defined by When measuring an observable quantity on a quantum system being in the state the probability of the outcome lying in an interval is where is the spectral measure of Consequently, a quantum state encapsulates a bunch of classical probability distributions, each corresponding to a physical quantity we are interested in.
 Let us single out a few observable quantities we are interested in, and let denote the random variable obtained by measuring in the initial state and let denote the random variable obtained by measuring in the final state According to (5), the squared OT distance of the quantum states should read as
 where the infimum is taken over all possible couplings of the quantum states and According to the convention introduced by De Palma and Trevisan [18], the set of all couplings of the quantum states is denoted by and is given by
-That is, a coupling of and is a state on such that
+where the transpose of a linear operator acting on is a linear operator on the dual space defined by the identity where and That is, a coupling of and is a state on such that
 and
-for all Note the clear analogy of the above definition of quantum couplings with the definition of classical couplings that can be rephrased as follows: is a coupling of and if
-and
+for all bounded Note the clear analogy of the above definition of quantum couplings with the definition of classical couplings that can be rephrased as follows: is a coupling of and if
 for every continuous and bounded function defined on We remark that is never empty, because the trivial coupling belongs to .
@@
 For more detail on the latter concept of quantum couplings, the interested reader should consult [9, 10, 27, 28, 32, 34, 30].
-Let us jump back to (6) and note that if the state of our composite quantum system is described by then
-Therefore, in view of (6), (7), and (9), the quadratic Wasserstein distance of and with respect to the measurable quantities is given by
-There is a correspondence between quantum couplings of the states (in the sense of (7)), and quantum channels (that is, completely positive, trace-preserving linear maps) defined on the support of that map to .
-This one-by-one relation is constructive and explicit, see Subsection 2.2.
-and especially Proposition 2 of [17] for further details.
-A consequence of this explicit relation — which we will heavily use in the sequel — is that the cost of a coupling can be expressed in terms of the corresponding channel as follows:
-See Subsection 5.2. of [18] for the justification of the above relation. Consequently, the quadratic quantum Wasserstein distances are obtained by an optimization over quantum channels:
-An important feature of the quadratic Wasserstein distances is that the distance of a state from itself (which may be positive) is always realized by the identity channel — see [17, Corollary 1].
-In addition to that, the following concavity-like result is true for any choice of and any :
+Let us jump back to (6) and note that by Born’s rule on quantum measurement, if the state of our composite quantum system is described by then
+Here we used the convention that the last line of (9) is defined to be if there is an eigenvector of outside the domain of for some Therefore, in view of (6), (7), and (9), the quadratic Wasserstein distance of and with respect to the measurable quantities is given by
+We recall (see [18, Definition 6]) that the energy of a state with respect to the observable is given by where is the spectral decomposition of and if for some The energy of with respect to the collection of observables is defined by Recall that classical quadratic Wasserstein spaces consist of probability measures with finite second moment.
+As the natural quantum analogs of them are states with finite energy, we restrict our attention to such quantum states in the sequel.
+By [18, Proposition 3], if the states have finite energy, then any quantum coupling has finite cost.
+Moreover, both and are trace-class for every – see [18, Lemma 3].
+Consequently, by the definition of Hilbert-Schmidt and trace-class operators, and are Hilbert-Schmidt, and so are and as taking the adjoint is an involution of Note furthermore that both and are Hilbert-Schmidt by definition, and hence the operators are trace-class as they are products of two Hilbert-Schmidt operators.
+A prominent coupling of a state with itself is the canonical purification which is the rank-one projection corresponding to the unit vector obtained from by the canonical isomorphism between and
+An important feature of the quadratic Wasserstein distances is that the distance of a state from itself (which may be positive) is always realized by the canonical purification — see [18, Corollary 1].
+That is,
+We have seen that the finite energy condition on implies that both and are Hilbert-Schmidt operators, and hence not only but also and and are trace-class operators.
+We note that taking the adjoint leaves both the Hilbert-Schmidt norm and the trace invariant, and hence and Consequently,
+Moreover, the following concavity-like result is true for any choice of and for any with finite energy:
 Indeed, (12) is an easy consequence of Theorem 1 and Corollary 1 of [18]. The following conjecture was popularized by Dario Trevisan and Giacomo De Palma in the fall of 2022 — see also [19].
@@
 A modified version of the quantum optimal transport distance (10) defined by
-is a true metric for all cost operator of the form
-— up to some non-degeneracy assumptions on the ’s generating to ensure the definiteness of that is, that only if
-The fact that the self-distance is realized by the identity channel for every [17, Corollary 1] and (12) together imply that
-We denote by the set of rank-one ortho-projections on that is, the set of pure states.
-Remark 1.
-If either or is pure then the only coupling is the tensor product and the only corresponding channel is , with
+is a true metric for all finite collections of observables on the set of those states on that have finite energy with respect to — up to some non-degeneracy assumptions on the ’s to ensure the definiteness of that is, that only if
+Let us denote by the set of rank-one ortho-projections on that is, the set of pure states.
+Note that if either or is a pure state, then the only coupling of them is the tensor product, that is, Therefore, by (10), the quadratic quantum Wasserstein distance has the following explicit form in this special case:
+Here we used that if and that for and that the operators are trace-class.
 2. Triangle inequality for quantum Wasserstein divergences — the proof of Theorem 1
@@
 Theorem 1.
+Let be a separable Hilbert space, and let be an arbitrary finite collection of observable quantities, and let be the corresponding quadratic quantum Wasserstein divergence defined by (13) and (10).
+Let and assume that or both and are in Moreover, assume that and have finite energy with respect to Then the triangle inequality
+holds true.
 Proof.
-If is pure or both and are pure, then by by Remark 1 and (16) we have
+If is pure or both and are pure, then by (14) and (11) we have
 and
@@
 By (12) the quantum Wasserstein divergence is a non-negative real number and hence eq. (15) is equivalent to
-If and are self-adjoint, and , then by the Cauchy-Schwarz inequality for the Hilbert-Schmidt inner product on and , and hence we get the following upper bound for the right-hand side (RHS) of (21):
+If is a state having finite energy with respect to the observable then is a Hilbert-Schmidt operator. Indeed, let be the spectral resolution of and let us compute the trace of the positive operator by
+Now let us note that this is precisely the trace of that we already have shown to be a trace-class operator. This latter statement can be checked by the direct computation
+So both and are Hilbert-Schmidt operators, and by the Cauchy-Schwarz inequality for the Hilbert-Schmidt inner product of them we get
+where we used that both and are trace-class and hence their traces coincide, and hence we get the following upper bound for the right-hand side (RHS) of (19):
 Now a Cauchy-Schwartz for the Euclidean space tells us that
-where we used again the Cauchy-Schwarz inequality for the Hilbert-Schmidt inner product for This completes the proof of (21) and hence that of the Theorem. ∎
+where we used again the Cauchy-Schwarz inequality for the Hilbert-Schmidt inner product
+for and This completes the proof of (19) and hence that of the Theorem.
 3. Numerical evidence for the triangle inequality for generic triplets of states
@@
 These minima are convincing but do not tell too much about how the gap depends on the states involved. Therefore, we worked out the following illustrative examples.
-In the first example, we considered and chose the deterministic states and We singled out the section of the Bloch ball, that is, we took ’s of the form
+In the first example, we considered and chose the deterministic states
+We singled out the section of the Bloch ball, that is, we took ’s of the form
 where The set of self-adjoint matrices generating the quadratic cost operator is chosen to be We considered the gap
@@
 The positivity condition is equivalent to the Euclidean length of being at most
-In this section we stick to and we choose the cost operator (15) to be
+In this section we stick to and we choose the cost operator to be
 where the “s” in the subscript of refers to the symmetry of meaning that it involves all the Pauli matrices.
@@
 we deduce that the cost function can be estimated from below by the Hellinger distance of and that is,
+5. Applications
+The aim of this section is to highlight a direct application of our main result in quantum complexity theory (Subsection 5.1), and to present a brief description of the role of quantum optimal transport distances in mathematical physics (Subsections 5.2 and 5.3).
+5.1. Wasserstein complexity
+We may define the Wasserstein complexity of quantum channels relying on quantum Wasserstein divergences.
+Given a finite collections of observables , let us define the Wasserstein complexity of a channel by
+In [39] the quantum Wasserstein complexity is introduced in a similar way, using the quantum generalization of the Hamming-Wasserstein classical metric.
+A connection between this complexity and the circuit cost of unitary channels is provided, the latter being bounded from below by the former.
+The argument of this bound relies heavily on the triangle inequality of the Wasserstein distance used.
+We suggest now that (37) may have a similar application in light of the triangle inequality proposed in this paper.
+To justify our suggestion we now prove a few useful porperties of (37) that is desirable of a complexity quantity.
+It follows from the definition, the positive definity of and the lack of self-distance in terms of that the Wasserstein complexity is faithful: if and only if is the identity.
+The Wasserstein complexity is subbadditive under concatenation: .
+Indeed,
+where the first inequality follows from the triangle inequality for the Wassertein divergence , the second inequality is due maximizing terms of a sum separately, and the last inequality follows from broadening the domain of the second maximum.
+From this it also follows directly that the Wasserstein complexity is subadditive under tensor products in the following sense:
+We note that whether the Wassertein complexity (37) is convex or not is an open question.
+5.2. Mean-field approximations of evolution equations
+Among the many applications of classical optimal transport, one is particularly important for the development of quantum optimal transport theory.
+In [20], Dobrushin employed a special approximation based on a transport-related metric, the so-called Kantorovich-Rubinstein metric, to prove the uniqueness of the solution to the Vlasov equation.
+The Vlasov equations, which describe the limiting situation of weakly interacting particles with a large radius of interaction, are among the most frequently used kinetic equations in statistical mechanics.
+More recently, Golse, Paul, and Mouhot have extended Dobrushin’s approach to the quantum setting.
+Since Dobrushin used an optimal transport metric to compare -particle densities and their mean-field limits, it was a natural idea to define a transport-related quantity for the purpose of comparing quantum states [27].
+The significance of the quantum counterpart of optimal transport (and the Wasserstein metric) in addressing various problems in quantum dynamics became even more evident later through their subsequent papers, see e.g.
+[9, 10, 27, 28, 31, 33].
+5.3. The Luttinger model
+The non-equilibrium dymamics of the Luttinger model, describing the low energy physics in Luttinger liquid is a field studied extensively.
+The question of how much the time evolved state described by differs from the initial state is often investigated.
+The paper [3] study the time evolution of Uhlmann fidelity (or the Loschmidt echo) which measures the overlap between the time evolved and the initial thermal equilibrium states, is evaluated for arbitrary initial temperatures and quench protocols.
+Instead of fidelity we can consider other dissimilarity measures between states to describe these effects.
+With zero initial temperature, i.e.
+the initial state is the pure ground state of the Hamiltonian, the Wasserstein distance is easiliy computable and can be used for further investigations.
 6. Appendix: the code used for the numerical study